# LoopServe: An Adaptive Dual-phase LLM Inference Acceleration System for Multi-Turn Dialogues

### v1

@@ 1. Introduction
  • We propose a new benchmark with 11 diverse multi-turn datasets. This benchmark incorporates diverse query positions and multi-turn dependencies, offering a comprehensive evaluation framework for LLMs in realistic long-context conversational tasks.
  • Extensive experiments conducted on 11 long-context multi-turn datasets demonstrate that LoopServe consistently outperforms state-of-the-art LLM acceleration models.
2. Preliminary and Related Work
In this section, we first introduce the preliminary of LLMs in Section 2.1, and then introduce various techniques that enhance the efficiency of long-context LLM inference in Section 2.2. Finally, we analyze existing benchmarks for LLM acceleration in Section 2.3.
2.1. Large Language Models
LLMs, such as GPT (Brown et al., 2020), Llama (Grattafiori et al., 2024), and DeepSeek (DeepSeek-AI, 2024, 2025), demonstrate exceptional abilities in context understanding and reasoning, achieving significant success in various tasks, such as data cleaning (Feuer et al., 2024; Sun et al., 2024; Ni et al., 2024; Zhu et al., 2024c; Eltabakh et al., 2024; Shankar et al., 2024; Ang et al., 2024; Yan et al., 2024), data management (Li et al., 2024e; Zhang et al., 2024; Halevy et al., 2024; Lou et al., 2024; Huang et al., 2024; Le et al., 2024; Tang et al., 2024; Nagrecha and Kumar, 2023), data visualization (Arora et al., 2023; Wu et al., 2024), data security (Li et al., 2024a), recommendation system (Yao et al., 2023; Zheng et al., 2024), text to SQL (van Renen et al., 2024; Fan et al., 2024; Li et al., 2024d; Zhu et al., 2024a; Ren et al., 2024; Luoma and Kumar, 2025; Zhang et al., 2023a), and information extraction (Zhu et al., 2024b; Xue et al., 2024; Singh et al., 2023; Hu et al., 2024; Tang et al., 2024; Dargahi Nobari and Rafiei, 2024; Chen and Koudas, 2024; Omar et al., 2025; Yu et al., 2025; Huo et al., 2024). Their impressive performance stems from extensive training datasets and a transformer architecture, which effectively captures long-range dependencies and enables advanced autoregressive generation.
2.1.1. Transformer Decoder
The effectiveness of Transformer decoders stems from the Multi-Head Self-Attention (MHSA) mechanism, which efficiently captures both local and global dependencies among input tokens. For clarity, we analyze a single representative Transformer block, as the operational principles are consistent across all blocks in the architecture. Given an input token sequence ${X}=[x_{1},x_{2},\cdots,x_{n}]$ with embeddings $\mathbf{X}\in\mathbb{R}^{n\times d}$, the self-attention mechanism computes weighted aggregations over the token sequence, where the weights are derived from inter-token similarity scores. Specifically, the MHSA mechanism generates query vectors $\mathbf{Q}^{i}\in\mathbb{R}^{n\times d_{k}}$, key vectors $\mathbf{K}^{i}\in\mathbb{R}^{n\times d_{k}}$, and value vectors $\mathbf{V}^{i}\in\mathbb{R}^{n\times d_{v}}$ for the $i$-th attention head through learned linear transformations:
(1) $\mathbf{Q}^{i}=\mathbf{X}\mathbf{W}_{Q^{i}},\quad\mathbf{K}^{i}=\mathbf{X}\mathbf{W}_{K^{i}},\quad\mathbf{V}^{i}=\mathbf{X}\mathbf{W}_{V^{i}},$
where $\mathbf{W}_{Q^{i}}\in\mathbb{R}^{d_{x}\times d_{k}}$, $\mathbf{W}_{K_{i}}\in\mathbb{R}^{d_{x}\times d_{k}}$ and $\mathbf{W}_{V_{i}}\in\mathbb{R}^{d_{x}\times d_{v}}$ are learnable parameters. The self-attention operation is subsequently applied to each triple $(\mathbf{Q}^{i},\mathbf{K}^{i},\mathbf{V}^{i})$, yielding the output of the $i$-th attention head $\mathbf{Z}^{i}$ as follows:
(2) $\mathbf{Z}^{i}=\textsf{Attention}(\mathbf{Q}^{i},\mathbf{K}^{i},\mathbf{V}^{i})=\textsf{Softmax}\left(\frac{\mathbf{Q}^{i}(\mathbf{K}^{i})^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}^{i},$
where the scaling factor $\sqrt{d_{k}}$ ensures numerical stability (Vaswani, 2017). To capture diverse semantic relationships, $h$ parallel attention heads process $\mathbf{X}$ concurrently. Then, $\mathbf{Z}=\textsf{Concat}(\mathbf{Z}^{1},\mathbf{Z}^{2},\dots,\mathbf{Z}^{h})\mathbf{W}_{O},$ where Concat is the concatenation operation, and $\mathbf{W}_{O}\in\mathbb{R}^{h\cdot d_{v}\times d_{o}}$ is the trainable projection matrix. Following the self-attention mechanism, the output is passed through a Feed Forward Network, whose result serves as the input to the next Transformer block.
2.1.2. Auto-regressive Generation Mechanism
LLMs utilize an autoregressive generation paradigm, where text is produced sequentially with each token conditioned on the tokens generated before it. This step-by-step approach maintains semantic coherence and ensures contextual relevance across the output. Formally, given an input sequence $X=[x_{1},x_{2},\cdots,x_{n}]$, LLM predicts the subsequent token $x_{n+1}$ by estimating the conditional probability distribution:
(3) $P(x_{n+1}|x_{1},x_{2},\cdots,x_{n})=\textsf{Softmax}(\mathbf{h}_{n}\mathbf{W}_{\text{out}}+\mathbf{b}_{\text{out}}),$
[…]

### v2

@@ 1 Introduction
  • We introduce a benchmark of 11 long-context multi-turn datasets with varied query positions and dependencies for realistic evaluation.
  • Experiments on 11 multi-turn datasets demonstrate the superior performance of LoopServe.
2 Preliminary and Related Work
2.1 Large Language Models
LLMs like GPT (Brown et al., 2020), Llama (Grattafiori et al., 2024), and DeepSeek (DeepSeek-AI, 2024; 2025) excel at context understanding and reasoning, enabled by large-scale training and the Transformer architecture (Vaswani, 2017). Transformers are effective due to Multi-Head Self-Attention (MHSA), which captures both local and global token dependencies. Given an input sequence $X=[x_{1},x_{2},\cdots,x_{n}]$ with embeddings $\mathbf{X}\in\mathbb{R}^{n\times d}$, the MHSA computes query vectors $\mathbf{Q}^{i}\in\mathbb{R}^{n\times d_{k}}$, key vectors $\mathbf{K}^{i}\in\mathbb{R}^{n\times d_{k}}$, and value vectors $\mathbf{V}^{i}\in\mathbb{R}^{n\times d_{v}}$ for the $i$-th attention head as $\mathbf{Q}^{i}=\mathbf{X}\mathbf{W}_{Q^{i}},\mathbf{K}^{i}=\mathbf{X}\mathbf{W}_{K^{i}},\mathbf{V}^{i}=\mathbf{X}\mathbf{W}_{V^{i}}$, where $\mathbf{W}_{Q^{i}}$, $\mathbf{W}_{K^{i}}$, and $\mathbf{W}_{V^{i}}$ are learnable matrices. Each $i$-th attention head $\mathbf{Z}^{i}$ as : $\mathbf{Z}^{i}=\textsf{Attention}(\mathbf{Q}^{i},\mathbf{K}^{i},\mathbf{V}^{i})=\textsf{Softmax}\left(\frac{\mathbf{Q}^{i}(\mathbf{K}^{i})^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}^{i}.$ Next, outputs from $h$ heads are concatenated and projected: $\mathbf{Z}=\textsf{Concat}(\mathbf{Z}^{1},\mathbf{Z}^{2},\dots,\mathbf{Z}^{h})\mathbf{W}_{O},$ where $\mathbf{W}_{O}$ is a learned projection. For text generation, LLMs use an autoregressive process: given $X=[x_{1},\ldots,x_{n}]$, the model predicts the next token $x_{n+1}$ by modeling $P(x_{n+1}|x_{1},x_{2},\cdots,x_{n})=\textsf{Softmax}(\mathbf{h}_{n}\mathbf{W}_{\text{out}}+\mathbf{b}_{\text{out}})$, where $\mathbf{h}_{n}$ is the state at step $n$. The next token $x_{n+1}$ is sampled from this distribution and appended to the sequence. Generation continues until an end-of-sequence token or a maximum length is reached.
2.2 Efficient Long-Context Inference
The performance of LLMs degrades with long input contexts, due to the quadratic complexity of self-attention, which scales as $O(Lhn^{2})$ for $L$ layers, $h$ heads, and sequence length $n$ (Li et al., 2024a). It makes long-sequence processing prohibitively expensive.
Context Compression Methods. Context compression methods reduce the effective sequence length, transforming lengthy inputs into more manageable representations. Filtering-based approaches such as LLMLingua (Jiang et al., 2023), LLMLingua-v2 (Pan et al., 2024), and CompAct (Yoon et al., 2024) focus on identifying and preserving high-relevance content, allowing models to process only critical information. In contrast, RAG-based (retrieval-augmented generation) methods (Zhao et al., 2024; Jiang et al., 2024b; Wang et al., 2025; Chen et al., 2025; Edge et al., 2024) construct knowledge graphs or extract semantic triples from the input, synthesizing them into condensed forms for LLMs. These strategies substantially decrease computational and memory costs, but may sacrifice fine-grained details, potentially affecting output quality.
Also, to reduce computational burden, KV-based approaches minimize the number of attention weight calculations during both prefilling and decoding, summarized in Table 3 in the Appendix.
Prefilling-stage Optimization. Self-attention requires $O(n^{2})$ computation for an input of length $n$. Recent methods such as Minference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and CritiPrefill (Lv et al., 2024) use binary masks $\mathbf{M}\in\{0,1\}^{n\times n}$ to zero out less important attention weights: $\min\,\left|\mathbf{{A}}^{j}_{i}-\mathbf{\hat{A}}^{j}_{i}\right|,s.t.\ ,\mathbf{\hat{A}}^{j}_{i}=\textsf{Softmax}\left(\frac{\mathbf{Q}^{j}(\mathbf{K}^{j})^{\top}}{\sqrt{d_{k}}}-c(1-\mathbf{M})\right)$, where $c$ is a large constant that suppresses masked entries. This reduces complexity to $O(\alpha n^{2})$ per head, with $\alpha\ll 1$. However, Minference (Jiang et al., 2024a) and CritiPrefill (Lv et al., 2024) rely on fixed attention patterns or block selection, while FlexPrefill (Lai et al., 2025) adjusts sparsity globally with simple heuristics. However, our experiments (Section 3) show that attention patterns are highly input-dependent and dynamic, so these static or coarse methods struggle to adapt in multi-turn scenarios.
Decoding-stage Optimization. During autoregressive generation, KV cache methods such as H2O (Zhang et al., 2023), SnapKV (Li et al., 2024b), AdaKV (Feng et al., 2024), and others (Ge et al., 2024; Li et al., 2024b) select important tokens to store, reducing redundancy. These approaches assume that critical tokens are near the end of the input, performing well on benchmarks like LongBench (Bai et al., 2024a; b), where queries are always placed last. However, as analyzed in Section 3.2, their effectiveness drops when queries appear elsewhere, underscoring the need for adaptive, context-aware KV selection in real dialogue.
3 Motivational Experiments and Insights
To clarify the core challenges in accelerating LLM inference for multi-turn dialogues, we conduct motivational experiments in Section 3.1 and 3.2. They reveal how attention patterns are dynamically sparse and how query position influences acceleration effectiveness. Building on these findings, we introduce the LoopServe system, specifically designed to address these real-world challenges.
3.1 Key Point 1: Uncertain Attention Patterns
As investigated previously, attention head matrices are highly sparse Jiang et al. (2024a). Existing acceleration methods (Xiao et al., 2024; Feng et al., 2024; LI et al., 2025; Li et al., 2024b) often rely on sparsifying attention matrices or selecting important KV tokens based on the assumption that attention patterns are fixed and can be identified offline. In reality, these patterns are highly variable across inputs, heads, and layers, limiting the effectiveness of such approaches, as reveled as follows.
(a) Vertical and slash lines.
(b) All heads of Llama 3.1.
(c) Average ratio on all heads.
(d) Overlap among different inputs.
Figure 1: Attention head sparsity is shown in (b) and (c).
Motivational Observation 1: Only 10% of vertical and slash lines can collectively account for most (e.g., 90%) of the attention weight. As shown in Figure 1 (a), in an attention matrix, a vertical line (column) indicates a single token attended by all others, common for special tokens like separators or keywords. A slash line (diagonal) shows each token mostly focusing on its nearby tokens, reflecting local attention patterns. We analyze this using the SAMSum QA dataset (Bai et al., 2024a) and the Llama-3.1-8B-Instruct with $n_{h}$ heads. For each query $X_{i}$ of length $n_{i}$, each $k$-th attention matrix $\mathbf{A}^{k}_{i}$ contains $n_{i}$ vertical lines $\mathcal{V}^{k}_{i}$ and $n_{i}$ slash lines $\mathcal{S}^{k}_{i}$. The total attention weight for a slash line $s^{k}$ is $\sum_{(a,b)\in s^{k}}\mathbf{A}^{k}_{i}[a][b]$, and for a vertical line $v^{k}$, it is $\sum_{(a,b)\in v^{k}}\mathbf{A}^{k}_{i}[a][b]$. We select the top $\eta\cdot 2n_{i}$ slash and vertical lines ($\mathcal{\hat{S}}^{k}_{i}$, $\mathcal{\hat{V}}^{k}_{i}$) based on their total weights. For each head $k$, we compute the ratio of weight within these lines: $r^{k}_{i}=\frac{1}{|\mathcal{D}|}\sum_{X_{i}\in\mathcal{D}}\frac{\sum_{(a,b)\in\mathcal{\hat{S}}^{k}_{i}\cup\mathcal{\hat{V}}^{k}_{i}}\mathbf{A}^{k}_{i}[a][b]}{n_{i}}.$ Averaging over all $n_{h}$ heads yields the mean ratio. As shown in Figure 1 (b), higher $\eta$ increases cumulative attention weight. Figure 1 (c) shows that for both Llama and Mistral, just 10% of slash and vertical lines account for 90% of total attention, indicating highly concentrated attention and enabling efficient selection by focusing on these sparse lines.
Motivational Observation 2: The positions of the top vertical and slash lines within the same head vary across different user inputs. For a model $M_{\theta}$ with $n_{h}$ attention heads and dataset $\mathcal{D}={X_{i}}$, we select the top $\eta\cdot 2n_{i}$ vertical and slash lines ($\mathcal{\hat{V}}^{k}_{i}$, $\mathcal{\hat{S}}^{k}_{i}$) for each input $X_{i}$ and attention head $k$. For any pair of queries $X_{i}$, $X_{j}$, the overlap of their selected lines under head $k$ is: $r^{k}_{i,j}=\frac{|\mathcal{\hat{S}}_{i}^{k}\cap\mathcal{\hat{S}}_{j}^{k}|+|\mathcal{\hat{V}}_{i}^{k}\cap\mathcal{\hat{V}}_{j}^{k}|}{|\mathcal{\hat{S}}_{i}^{k}\cup\mathcal{\hat{S}}_{j}^{k}|+|\mathcal{\hat{V}}_{i}^{k}\cup\mathcal{\hat{V}}_{j}^{k}|}.$ Averaging over all heads and input pairs gives the mean overlap ratio: $\frac{1}{n_{h}|\mathcal{D}|^{2}}\sum_{k=1}^{n_{h}}\sum_{X_{i},X_{j}\in\mathcal{D}}r^{k}_{i,j}$. Using the SAMSum QA dataset (Bai et al., 2024a) and models Llama-3.1-8B-Instruct and Mistral-7B-Instruct-v0.3, with $\eta$ ranging from 0.1 to 0.3, Figure 1 (d) and Figure 7 in Appendix show that for most heads, the overlap remains below 0.5. This indicates that the most important lines differ significantly depending on the input, even within the same head. As a result, important lines cannot be reliably determined offline for use during online inference.
Motivational Observation 3: For an input $X_{i}=[C_{i}^{1},C^{2}_{i}]$ split into two segments, the top vertical and slash lines within the same head differ between $C^{1}_{i}$ and $C^{2}_{i}$. Each segment shows its own local attention sparsity pattern. As illustrated in Figure 2 (a), the key vertical and slash lines in $C^{1}_{i}$’s attention matrix are largely absent in $C^{2}_{i}$, which displays distinct local patterns. To verify this, we use the SAMSum QA (Bai et al., 2024a) dataset and Llama-3.1-8B-Instruct. For each $X_{i}$ in dataset $\mathcal{D}$, we split it into $C^{1}_{i},C^{2}_{i}]$ and extract the attention matrices $\mathbf{A}^{k}_{C^{1}_{i}}$ and $\mathbf{A}^{k}_{C^{2}_{i}}$ for each head $k$. After selecting the top-$\eta$ important slash and vertical lines for each ($L^{k}_{C^{1}_{i}}$, $L^{k}_{C^{2}_{i}}$) for ( $\mathbf{A}^{k}_{C^{1}_{i}}$, $\mathbf{A}^{k}_{C^{2}_{i}}$), we compute the overlap rate: $r^{k}_{C_{i}^{1}\to C_{i}^{2}}={\sum_{l\in L^{k}_{C^{1}}}\mathbb{I}(l\in L^{k}_{C_{i}^{2}})}/{|L^{k}_{C_{i}^{2}}|},$ where $\mathbb{I}$ indicates whether a line from $C^{1}_{i}$ is also important in $C^{2}_{i}$. Averaging across data gives the mean overlap line rate for each head. Figure 2 (b) shows that, for different $\eta$, the overlap in important lines between $C^{1}_{i}$ and $C^{2}_{i}$ is consistently low and unstable. This confirms that each segment exhibits unique local attention patterns. This finding indicates that using only a segment (such as the last window or last few tokens) to predict important attention patterns for the whole input is unreliable. As a result, acceleration methods like Minference (Jiang et al., 2024a), SnapKV (Li et al., 2024b), H2O (Zhang et al., 2023), and Keyformer (Adnan et al., 2024), which rely on such assumptions, struggle to deliver consistent performance in real-world scenarios.
3.2 Key Point 2: Question Position Matters.
Motivational experiments indicate that both prefilling based methods and decoding phase acceleration methods, which depend on offline sparse pattern discovery or fixed sparse patterns, tend to underperform in practical scenarios. However, their reported outcomes on existing benchmarks are often similar to those of large language models that use full attention. Why does this discrepancy occur? The main reason is that benchmarks like Longbench (Bai et al., 2024a; b) always place the user question $q_{i}$ at the end of the input $X_{i}=[C_{i},q_{i}]$, so the LLM answers $q_{i}$ based on context $C_{i}$. In this setup, acceleration methods only need to focus on the last observation window (near the question), which makes it easier to identify context tokens relevant to the question, thus partially mitigating the unpredictability found in real-world input patterns.
(a) Attention map.
(b) Overlap of different segments.
(c) Important token overlap.
(d) Question in different position.
Figure 2: Different attention sparsity patterns and query position impact on performance.
Motivational Observation 4: Relying only on the last observation window cannot reliably identify important input tokens for generating the output. Recent methods like H2O (Zhang et al., 2023), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024) select the top-$B$ important tokens $\hat{X}_{i}$ from $X_{i}$ based on attention between each input token and the last observation window $X^{obs}_{i}$ (the last $n_{s}$ tokens), formally as $\hat{X}_{i}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{j=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in X^{obs}_{i}}\mathbf{A}^{k}_{i}[a][b]$. However, the true top-$B$ tokens $\hat{X}^{*}_{i}$ for generating the output $Y_{i}$ should be selected based on their attention to the output tokens: $\hat{X}^{*}_{i}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{j=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in Y_{i}}\mathbf{A}^{k}_{i}[a][b]$. We measure the overlap $r_{i}$ between $\hat{X}_{i}$ and $\hat{X}^{*}_{i}$ ($B$ is set to 10% of $|X_{i}|$), using Llama-3.1-8B-Instruct and LongEval’s topic retrieval set, where the instruction $q_{i}$ is placed at the beginning, middle, or end of $C_{i}$. As shown in Figure 2 (c), the average overlap of important tokens is highest when the question is at the end, but much lower when it appears in middle or the beginning. This demonstrates that focusing only on the last part of the input misses relevant information unless the question is placed last.
Similarly, as shown in Figure 2 (d), SnapKV and AdaKV match the original model only when the question is at the end. Their performance drops sharply when the question appears earlier, since they rely on the last tokens for context selection. This shows that current methods are overly dependent on input order and do not generalize well when question positions vary.
4 Multi-turn Long-Context Benchmarks
Existing benchmarks, such as NumericBench (Li et al., 2025), LongBench (Bai et al., 2024a; b), and LongEval (Dacheng Li* & Zhang, 2023) focus on single-turn tasks and place user queries only at the context end, which do not reflect the complexity of real-world, multi-turn conversations (LI et al., 2025). We introduce a benchmark of 11 long-context multi-turn datasets with varied query positions and dependencies. Specifically, each $m$-turn instance is defined as $\mathcal{D}={\mathcal{I}_{i}=[(C_{i,1},q_{i,1},a_{i,1}),\dots,(C_{i,m},q_{i,m},a_{i,m})]}_{i=1}^{|\mathcal{D}|}$, where $C_{i,j}$ is the context at turn $j$ (possibly empty), $q_{i,j}$ is the user query, and $a_{i,j}$ is the LLM-generated answer. We ensure diversity by: (1) Query Position: For each turn, the query $q_{i,j}$ can appear at the beginning, end, or between paragraphs of $C_{i,j}$, reflecting more realistic query placements. (2) Query Relevance: Answers $a_{i,j}$ may depend on any subset of current and previous contexts $\{C_{i,1},\ldots,C_{i,j}\}$, with variable subset sizes, simulating diverse real-world dependencies. We construct multi-turn benchmarks for several tasks, including question answering, summarization, and few-shot learning. For construction procedures and detailed benchmark statistics, please refer to Appendix A.5.
5 LoopServe System
Figure 3: Framework overview of LoopServe.
As shown in Figure 3, we propose LoopServe, an adaptive dual-phase system that performs online attention sparsification during the prefilling phase and progressive KV compression during the decoding phase. For an $m$-turn input $\mathcal{I}_{i}=\{X_{i,j}\}_{j=1}^{m}$, where $X_{i,j}$ is the context or query in the $j$-th turn, LoopServe generates each answer $y_{i,j}$ using these two steps.
Step 1. Online Attention Head Sparsification in Prefilling. In Algorithm 1 (line 3-5) in Appendix, for each new input $X_{i,j}$ and all inputs $X_{i}=\left(\cup_{j^{\prime}=1}^{j-1}{(X_{i,j^{\prime}}\cup y_{i,j^{\prime}})}\right)\cup X_{i,j}$, we get $\hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$ as the new appended input in the $j$ turn. Then, for each $k$-th attention head, we first compute the attention matrix $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]\in\mathbb{R}^{\hat{n}_{i,j}\times n_{i}}$, and then select the slash lines $\mathcal{\hat{S}}^{k}_{i,j}$ and vertical lines $\mathcal{\hat{V}}^{k}_{i,j}$ that collectively recover at least $\alpha$ of the total attention weight in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$.
Step 2. Progressive KV Compression in Decoding. As described in Algorithm 1 (lines 7–14) in Appendix, after every re-selection interval $n_{d}$ tokens, the framework uses the ProgressiveSelection Algorithm 3 to compute a subset of input tokens $\hat{X}^{k}_{i}\subseteq X_{i}$ for each attention head $k$. these selected tokens $\hat{X}^{k}_{i}$ are important for output generation. At each decoding step, LoopServe leverages the compressed KV cache $\{\mathcal{\hat{S}}^{k}_{i,j^{\prime}},\mathcal{\hat{V}}^{k}_{i,j^{\prime}}\}_{j^{\prime}=1,k}^{j,n_{h}}$ to generate the output sequence $y_{i,j}$.
5.1 Online Attention Sparsification in Prefilling
As shown in Key Point 1 in Section 3.1, attention sparsity patterns are highly dynamic, making static or offline selection ineffective. To address this, we propose an online adaptive algorithm that, during prefilling, selects a subset of slash and vertical lines to recover at least an $\alpha$ fraction of the total attention weight for each head, which can be reused in later dialogue turns.
Definition 1 (Online Prefilling Sparsification Problem).
Given the LLM model $M_{\theta}$ with $n_{h}$ attention heads, the input $X_{i}=\left(\bigcup_{j^{\prime}=1}^{j-1}(X_{i,j^{\prime}}\cup y_{i,j^{\prime}})\right)\cup X_{i,j}$, where $y_{i,j^{\prime}}$ is the answer for $X_{i,j^{\prime}}$ and $X_{i,j}$ is the current turn’s input. We denote the concatenation of the previous answer $y_{i,j-1}$ and the current user input $X_{i,j}$ as $\hat{X}_{i,j}=[y_{i,j-1},X_{i,j}]$, whose corresponding attention matrix requires sparsification. The $k$-th attention matrix between $X_{i}$ and $\hat{X}_{i,j}$ is denoted as $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]\in\mathbb{R}^{\hat{n}_{i,j}\times n_{i}}$. Let $\mathcal{S}^{k}_{i,j}$ (resp., $\mathcal{V}^{k}_{i,j}$ ) denote the set of all slash lines (resp., vertical lines) in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$. The goal is to select a subset of slash lines $\hat{\mathcal{S}}^{k}_{i,j}\subseteq\mathcal{S}^{k}_{i,j}$ and a subset of vertical lines $\hat{\mathcal{V}}^{k}_{i,j}\subseteq\mathcal{V}^{k}_{i,j}$ such that together they recover at least an $\alpha$ fraction of the total attention weight in $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$, where $\alpha\in[0,1]$.
$\min\sum_{s\in\mathcal{\hat{S}}^{k}_{i,j}}l_{s}+\sum_{v\in\mathcal{\hat{V}}^{k}_{i,j}}l_{v},\quad s.t.\sum_{(a,b)\in(\mathcal{\hat{S}}^{k}_{i,j}\cup\mathcal{\hat{V}}^{k}_{i,j})}\mathbf{A}^{k}_{i}[a][b]\geq\alpha\cdot\hat{n}_{i,j},$ (1)
where $\hat{n}_{i,j}$ is the total attention weight of the matrix $\mathbf{A}^{k}_{i}[\hat{X}_{i,j}]$, and $l_{s}$ (resp., $l_{v}$) is the length of the slash line $s$ (resp., vertical line $v$).
Theorem 1.
The prefilling sparsification problem is NP-hard. The proof is detailed in Appendix A.6.
Algorithm. Algorithm 2 in Appendix A.7 takes as input the concatenated sequence $\hat{X}{i,j}=[y_{i,j-1},X_{i,j}]$, where $y_{i,j-1}$ is the previous answer and $X_{i,j}$ is the current user input. It also requires the $k$-th attention head of the LLM $M_{\theta}$ and a sparsity threshold parameter $\alpha$. The output consists of the selected slash lines $\mathcal{\hat{S}}^{k}_{i,j}$ and selected vertical lines $\mathcal{\hat{V}}^{k}_{i,j}$. The algorithm begins by sampling a subset $\tilde{X}_{i,j}$ from the concatenated input $\hat{X}_{i,j}$ to reduce computational cost. It then computes the query matrix $\mathbf{\tilde{Q}}^{k}_{i,j}$ for $\tilde{X}_{i,j}$ and the key matrix $\mathbf{K}^{k}_{i}$ for the full input $X_{i}$. Next, all slash lines $\mathcal{S}^{k}{i,j}$ and vertical lines $\mathcal{V}^{k}_{i,j}$ are summarized based on $\mathbf{A}^{k}_{i}[\tilde{X}{i,j}]$ and sorted in descending order. Two empty sets, $\mathcal{\hat{S}}^{k}_{i,j}$ and $\mathcal{\hat{V}}^{k}_{i,j}$, are initialized to store the selected lines, while the overlap weights $ol_{s}$ and $ol_{v}$ are initialized to zero. Algorithm 2 then iteratively selects lines until the total recovered attention weight sum meets or exceeds $\alpha\cdot\textsf{sum}(\mathbf{A}^{k}_{i}[\tilde{X}_{i,j}])$. At each iteration, it compares the top slash line $s\in\mathcal{S}^{k}_{i,j}$ and the top vertical line $v\in\mathcal{V}^{k}_{i,j}$ based on their marginal contributions, $\Delta w_{s}=w_{s}-ol_{v}$ and $\Delta w_{v}=w_{v}-ol_{s}$. Since each slash line overlaps with only one vertical line, $\Delta w_{s}=w_{s}-ol_{v}$. The line with the greater marginal contribution is added to its respective set ($\mathcal{\hat{S}}^{k}_{i,j}$ or $\mathcal{\hat{V}}^{k}_{i,j}$), and the overlap weights and total recovered weight are updated accordingly. The loop terminates once the recovery condition is satisfied, and the algorithm returns the sets $\mathcal{\hat{S}}^{k}_{i,j}$ and $\mathcal{\hat{V}}^{k}_{i,j}$. The time complexity is $O(n_{i}|\tilde{X}_{i,j}|+n_{i}\log n_{i}+n_{i})$, as detailed in Appendix A.7.1.
5.2 Progressive KV Compression in Decoding
As shown in Section 3.2, fixed or last-token-focused decoding struggles when queries are not at the end of the input. To overcome this, we propose a progressive KV compression that selects important input tokens based on recent outputs, leading to greater overlap with truly important input tokens.
To verify this, we consider the LLM $M_{\theta}$ with $n_{h}$ attention heads, an input sequence $X_{i}$, a generated output sequence $Y_{i}$, and a window size $n_{w}$. As in Section 3.2, we compute the ground-truth top-$B$ important tokens $\hat{X}^{*}_{i}\subseteq X_{i}$ using the output $Y_{i}$. We then use observation windows, extracted from either $X_{i}$ or $Y_{i}$, to select the top-$B$ important tokens. Specifically, given the window size $n_{w}$, the $\frac{|X_{i}|}{n_{w}}-$th to 1-1 observation windows from $X_{i}$ are defined as $X_{i}[n_{i}-m\cdot n_{w}:n_{i}-(m-1)\cdot n_{w}]$, $X_{i}[n_{i}-(m-1)\cdot n_{w}:n_{i}-(m-2)\cdot n_{w}]$, …, $X_{i}[n_{i}-n_{w}:n_{i}]$. The 0-th to $\frac{|Y_{i}|}{n_{w}}$ observation windows from $Y_{i}$ are $Y_{i}[0:n_{w}]$, $Y_{i}[n_{w}:2n_{w}]$, …, $Y_{i}[(m-1)\cdot n_{w}:m\cdot n_{w}]$. For each observation window $X^{obs}_{i}$ from $X_{i}$ or $Y_{i}$, we select the top-$B$ tokens $\hat{X}_{i}\subseteq X_{i}$ following Section 3.2, and compute the overlap rate between $\hat{X}_{i}$ and the ground truth $\hat{X}^{*}_{i}$ as $\frac{|\hat{X}_{i}\cap\hat{X}^{*}_{i}|}{B}$.
(a) Important token overlapping.
(b) Decoding blocks.
Figure 4: Progressive decoding.
We use LongEval (Dacheng Li* & Zhang, 2023) as in Section 3, experimenting with Llama-3.1-8B-Instruct and setting $B\in\{5\%,10\%,15\%\}\cdot|X_{i}|$. Figure 4(a) shows that important tokens selected from input blocks (block index <0<0) have low overlap rates, while those selected from output blocks ($\geq 0$) have much higher overlap. This indicates that using output tokens is more effective for identifying relevant input tokens.
We also observe that the overlap of important input tokens between output blocks decreases as the distance between blocks increases, which indicates that we can use the most recent output tokens to select important input tokens for the next output block. Specifically, given the output sequence $Y_{i}$, we divide it into $n_{b}$ blocks, i.e., $Y_{i}=[Y^{1}_{i},\cdots,Y^{n_{b}}_{i}]$, where $Y^{j}_{i}$ denotes the $j$-th output block of size $\frac{|Y_{i}|}{n_{b}}$. For each block $Y^{j}_{i}$, we compute the top-$B$ important input tokens as: $\hat{X}_{i,Y^{j}_{i}}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{k=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in Y^{j}_{i}}\mathbf{A}^{k}_{i}[a][b]$. Next, we compare the overlap $\frac{|\hat{X}_{i,Y^{j}_{i}}\cap\hat{X}_{i,Y^{j^{\prime}}_{i}}|}{B}$ of important input tokens between every pair of output blocks $Y^{j}_{i}$ and $Y^{j^{\prime}}_{i}$. As shown in Figure 4 (b), for each block $Y^{j}_{i}$, the overlap of important input tokens between $Y^{j}_{i}$ and $Y^{j^{\prime}}_{i}$ (where $j^{\prime}>j$) gradually decreases as the block distance increases. Notably, the overlap between $Y^{j}_{i}$ and its immediate successor $Y^{j+1}_{i}$ is higher compared to earlier blocks such as $Y^{j-2}_{i}$. This indicates that tokens identified from $Y^{j}_{i}$ are highly relevant for the generation of $Y^{j+1}_{i}$. Therefore, when generating $Y^{j+1}_{i}$, we can use the preceding block $Y^{j}_{i}$ to dynamically identify the top-$B$ important input tokens. Based on this, we propose the following progressive KV compression algorithm.
Algorithm. Algorithm 3 firstly set the answer $y_{i,j}$ as empty and decoding step counter $n_{o}$ to 0 (line 1). During the decoding loop, if the current step reaches either the empirically predefined decoding size 16 or a re-selection interval $n_{d}$ (line 3), the algorithm extracts the most recent tokens from the input sequence $X_{i}$, i.e., $X^{obs}_{i}=X_{i}[|X_{i}|-n_{d}:|X_{i}|]$), and dynamically updates the compressed KV cache subset $\hat{X}^{k}_{i}$ for each $k$-th attention head (line 4-5). Specifically, for each head, we select top-$B$ tokens for each head as: $\hat{X}^{k}_{i}=\arg\max_{\hat{X}^{k}_{i}\subseteq X_{i},|\hat{X}^{k}_{i}|=B|}\sum_{a\in\hat{X}_{i}}\sum_{b\in X^{obs}_{i}}\mathbf{A}^{k}_{i}[a][b],$ The LLM generates the next token using the compressed KV cache and the updated input, which is appended to both the input and output sequences (line 6-9). This process iterates until the LLM completes generation, returning the final answer $y_{i,j}$.
6 EXPERIMENTS
6.1 Experimental Settings
[…]